My cars cooling system holds 4.5 litres. I put 1.5 litres of pure antifreeze and 3 litres of water in it. (33.33% antifereeze?) I have since drained 1 litre of this mixture and topped up with 1 litre of pure antifreeze. Could you please advise me how much of this new mixture I need to drain in order to achieve a 50/50 mix of antifreeze/water by topping up with pure antifreeze?
Many many thanks in advance, Wayne

Answer by scott8148(6628)

AF starting amount ___ 1.5 or 3/2 L (or 9/6)

AF amount drained ___ (1.5)(1 / 4.5) = 1/3 L or 2/6

AF current amount ___ (9/6) - (2/6) + (6/6) = 13/6

current concentration ___ (13/6) / (9/2) = 13/27

[(9/2) - x](13/27) + (x)(100%) = (9/2)(50%)

(13/6) - (13x/27) + x = 9/4

multiplying by 108 (LCD) ___ 236 - 52x + 108x = 243 ___ 56x = 7

x = 1/8 L